The function l=-0.02x^2+60x models the income earned (i), in dollars, by selling units of product is:
What are the greatest possible income?
How many unit must be sold in order to have the greatest income?
How many unit must be sold to generte an income of $35,000?
HELP i have no clue how to do this!

Answer by stanbon(75887)

The function l=-0.02x^2+60x models the income earned (i), in dollars, by selling units of product is:
What are the greatest possible income?
The function is a quadratic with a=-0.02, b=60
The maximum value of the function occurs when x = -b/2a = -60/(2*-0.02) = 1500
The income when x=1500 is i = $45000.00
------------------------------------------------------
How many unit must be sold in order to have the greatest income?
Ans: 1500 units
---------------------
How many unit must be sold to generate an income of $35,000?
35000 = -0.02x^2+60x
0.02x^2-60x+35000 = 0
x^2 - 3000x + 175000 =0
x=59.51
Rounding up get x=60 units
===============================
Cheers,
Stan H.